OpenStudy (anonymous):

Plz help medal will be given giselle pays $210 in advance on her account at the athletic club.Ech time she uses the club,$15 is deducted from the account.Model the situation with a linear function. A.b = 210 + 15x B.b = 210 – 15x C.b = 195 + 15x D.b = 195 – 15x

OpenStudy (cj49):

she pays 210 advance and 15 is deducted each time from her accnt b=210+15x

OpenStudy (anonymous):

So I was wrong and it's A right?

OpenStudy (cj49):

wat did u choose

OpenStudy (anonymous):

b=210-15x

OpenStudy (cj49):

she paid 210 in advance and each time 15 wer deducted from her accnt...m not sure if that 15 is being deducted from that 210 or not.

OpenStudy (anonymous):

Not sure myself either.

OpenStudy (anonymous):

Well thanks for your help Ill just pick one and see if it's right later on

OpenStudy (cj49):

np.

OpenStudy (anonymous):

The correct answer is B=210-15x